OpenStudy (anonymous):

Write a possible explicit rule for the nth term of the sequence: 1, 1/2, 1/3, 1/4, 1/5, ... 1.an = 1/n-1 2.an = 1/an-1 3.an = 1/n2 4.an = 1/n

OpenStudy (anonymous):

1/n

OpenStudy (anonymous):

you sure?

OpenStudy (anonymous):

Let's check. since first term: n is 1, if you plug in 1 in the 1/n, you get one. second term: n=2, so, 1/n = 1/2. So my answer is correct It's good to check :)